Solve each rational equation. Complex numbers are very confusing for me.
3x over x+1
plus
4 over x-2
=
3
Thank you so much , I greatly appreciate the help

Answer by Boreal(15235)

3x/(x+1)+4/(x-2)=3
common denominator is (x+1)(x-2)
3x(x-2)+4(x+1)=3(x-2)(x+1)
3x^2-6x+4x+4=3(x^2-x-2)
3x^2-2x+4=3x^2-3x-6
-2x+4=-3x-6
x=-10
check
-30/-9 +4/-12=3
3 1/3 - (1/3)=3
checks